# Explore until Confident: Efficient Exploration for Embodied Question Answering

### v1

@@ II Related Work
VLMs for other embodied tasks. VLMs have shown promise in solving other embodied tasks using their commonsense reasoning and knowledge [wen2023road, driess2023palm, sumers2023distilling, shen2023distilled]. VLMs have been fine-tuned to exhibit strong physical and spatial reasoning [chen2024spatialvlm, gao2023physically, brohan2023rt]. The closest work to ours include [kwon2023toward], where queries to a VLM and subsequently an LLM are used to help the robot reason about the most informative angle from which to view a scene, e.g., the angle that helps determine whether a coffee cup is empty and needs to be cleaned up. Similarly [shah2023navigation] uses an LLM to generate search heuristics guiding robot finding objects in unknown environments. Recently visual prompting with VLMs has been proposed to acquire VLM’s prior knowledge about specific regions on 2D images [nasiriany2024pivot, liu2024moka], e.g., searching for directions for reaching an object. In our work, we go beyond simple active exploration strategies generated by LLMs/VLMs, and carefully consider uncertainty estimation and active exploration while using the commonsense reasoning capabilities of VLMs to determine where to navigate in larger scenes.
@@ V-B Applying Multi-Step CP for Embodied Question Answering
Next, we describe how CP provides a principled and more interpretable stopping criterion for multi-step exploration by building on the multi-step CP approach presented in [ren2023knowno]. In particular, we consider datapoints corresponding to episode-level sequences of inputs. By performing calibration at the sequence level using a carefully chosen non-conformity score function, we can ensure that prediction sets can be constructed causally (i.e., time-step by time-step) at test time.
@@
[…]
@@
[…]
@@
At test time, we thus construct the set $C^{t}(x^{t}_{\text{test}})$ at each step and maintain the intersection of these sets over time. If the resulting intersection contains only a single element, the robot halts its exploration with $1-\epsilon$ confidence that the corresponding answer is correct. Alternately, if the maximum allowable time horizon $T$ is reached and the intersected set contains multiple answers, or the intersected set is empty, the robot returns the answer $y$ with highest $\hat{f}_{y}(x^{t})$ from time $t$ with the highest $\text{Rel}(x^{t})$.
@@ VI HM-EQA Dataset
While prior work has primarily considered synthetic scenes and simple questions such as “what is the color of the coffee table?” or “how many sofas are there in the living room?” involving basic attributes of relatively large pieces of furniture, we are interested in applying our VLM-based framework in more realistic and diverse scenarios, where the question can be more open-ended and possibly require semantic reasoning. To this end, we propose HM-EQA, a new EQA dataset based on the Habitat-Matterport 3D Research Dataset (HM3D), which provides hundreds of photo-realistic, diverse indoor 3D scans [ramakrishnan2021habitat].
@@ VII-B Q1: Semantic Exploration - Baselines
For evaluation, we vary the number of time steps (normalized by $T_{\epsilon}$ of each scenario, defined in Section VI) allowed — the robot does not need to determine when to stop in these experiments (we will consider the different stopping criteria in Q2; Sec. VII-E). For all methods including ours, in each scenario the robot stops at the maximum number of time steps allowed, and the final answer is chosen to be $y$ with highest $\hat{f}_{y}(x^{t})$, at time $t$ with the highest $\text{Rel}(x^{t})$. Also, since we do not use conformal prediction as a stopping criterion for these experiments, we do not need to split the data into calibration and test scenarios. We can thus use all 500 scenarios from the dataset for evaluation.
@@ VII-D Q1: Semantic Exploration - Comparing to CLIP-FBE
VII-E Q2: Stopping Criterion - Baselines
@@
VII-F Q2: Stopping Criterion - Simulation Results
@@
VII-G Q2: Stopping Criterion - Hardware Results
@@ Acknowledgments
We thank Donovon Jackson, Derick Seale, and Tony Nguyen for contributing to the HM-EQA dataset. The authors were partially supported by the Toyota Research Institute (TRI), the NSF CAREER Award [#2044149], and the Office of Naval Research [N00014-23-1-2148]. This article solely reflects the opinions and conclusions of its authors and NSF, ONR, TRI or any other Toyota entity.

### v3

@@ II Related Work
VLMs for other embodied tasks. VLMs have shown promise in solving other embodied tasks using their commonsense reasoning and knowledge [wen2023road, driess2023palm, sumers2023distilling, shen2023distilled]. VLMs have been fine-tuned to exhibit strong physical and spatial reasoning [chen2024spatialvlm, gao2023physically, brohan2023rt]. The closest work to ours include [kwon2023toward], where queries to a VLM and subsequently an LLM are used to help the robot reason about the most informative angle from which to view a scene, e.g., the angle that helps determine whether a coffee cup is empty and needs to be cleaned up. Similarly [shah2023navigation] uses an LLM to generate search heuristics guiding robot finding objects in unknown environments. Recently visual prompting with VLMs has been proposed to acquire VLM’s prior knowledge about specific regions on 2D images [nasiriany2024pivot, liu2024moka], e.g., searching for directions for reaching an object. In our work, we go beyond simple active exploration strategies generated by LLMs/VLMs, and carefully consider uncertainty estimation and active exploration while using the commonsense reasoning capabilities of VLMs to determine where to navigate in larger scenes. Previous work [tian2022vibus] has also explored quantifying VLM uncertainty in semantic understanding, as well as leveraging VLM uncertainty for better anomaly detection in autonomous driving scenarios [tian2023unsupervised].
@@ V Stopping Criterion for Exploration and Answering the Question
Figure 4: For determining when to stop, we apply a principled approach based on multiple-step conformal prediction: at each step, a prediction set is generated, and the robot keeps the intersection of the sets until there is only one option remaining. The correct answer is guaranteed to be the remaining one with user-specified probability.
@@ V-B Applying Multi-Step CP for Embodied Question Answering
Next, we describe how CP provides a principled and more interpretable stopping criterion for multi-step exploration by building on the multi-step CP approach presented in [ren2023knowno]. In particular, we consider datapoints corresponding to episode-level sequences of inputs. By performing calibration at the sequence level using a carefully chosen non-conformity score function, we can ensure that prediction sets can be constructed causally (i.e., in the order of time-step) at test time.
@@
For all time $t\in[T]$, the causally constructed prediction set $C^{t}(x^{t}_{\text{test}})$ contains the sequence-level set $\bar{C}(\bar{x}_{\text{test}})$. Moreover, $\cap_{t=0}^{T}C^{t}(x^{t}_{\text{test}})=\bar{C}(\bar{x}_{\text{test}})$.
@@
With probability $1-\epsilon$ for test scenarios drawn from $\mathcal{D}$, the ground-truth label $y_{\text{test}}$ is contained in the prediction set $\cap_{k=0}^{t}C^{k}(x^{k}_{\text{test}})$ for all $t\in[T]$.
@@
At test time, we thus construct the set $C^{t}(x^{t}_{\text{test}})$ at each step and maintain the intersection of these sets over time. If the resulting intersection contains only a single element, the robot halts its exploration with $1-\epsilon$ confidence that the corresponding answer is correct (Fig. 4). Alternately, if the maximum allowable time horizon $T$ is reached and the intersected set contains multiple answers, or the intersected set is empty, the robot returns the answer $y$ with highest $\hat{f}_{y}(x^{t})$ from time $t$ with the highest $\text{Rel}(x^{t})$.
@@ VI HM-EQA Dataset
Figure 5: Sample scenarios from the HM-EQA dataset. The images are the views used by the robot to determine the final answer in our experiments. The boxes show the questions with the true answers bolded.
While prior work has primarily considered synthetic scenes and simple questions such as “what is the color of the coffee table?” or “how many sofas are there in the living room?” involving basic attributes of relatively large pieces of furniture, we are interested in applying our VLM-based framework in more realistic and diverse scenarios, where the question can be more open-ended and possibly require semantic reasoning. To this end, we propose HM-EQA, a new EQA dataset based on the Habitat-Matterport 3D Research Dataset (HM3D), which provides hundreds of photo-realistic, diverse indoor 3D scans [ramakrishnan2021habitat]. Sample scenes are shown in Fig. 5.
@@
VI-A Extension: Answering Questions without Multiple Choices
While the HM-EQA dataset contains multiple-choice questions instead of ones without multiple choices and we believe this is a reasonable design choice for benchmarking purposes, our framework can be also be extended to handle questions without multiple choice answers. We can apply the multiple-choice question answering (MCQA) setup in [ren2023knowno] where an LLM (VLM in our case) first proposes different possible answers given the input (image observation and question in our case) based on few-shot examples, and then chooses one among them. To ensure coverage of the ground truth answer, we will add an additional option ‘An option not listed here’ in case the VLM does not generate the correct one. This setup allows us to quantify VLM’s uncertainty among different possible answers in Section V-B. Fig. 6 shows an example of GPT4-V generating possible answers to a question given the view.
Figure 6: Prompt (yellow) and output (blue) of few-shot prompting GPT4-V to generate possible answers to the question given the view. The correct answer is ‘C) Cartoon network’. The actual prompt used has images encoded into a text form.
@@ VII-B Q1: Semantic Exploration - Baselines
For evaluation, we vary the number of time steps (normalized by $T_{\epsilon}$ of each scenario, defined in Section VI) allowed — the robot does not need to determine when to stop in these experiments (we will consider the different stopping criteria in Q2; Sec. VII-F). For all methods including ours, in each scenario the robot stops at the maximum number of time steps allowed, and the final answer is chosen to be $y$ with highest $\hat{f}_{y}(x^{t})$, at time $t$ with the highest $\text{Rel}(x^{t})$. Also, since we do not use conformal prediction as a stopping criterion for these experiments, we do not need to split the data into calibration and test scenarios. We can thus use all 500 scenarios from the dataset for evaluation.
@@ VII-D Q1: Semantic Exploration - Comparing to CLIP-FBE
VII-E Q1: Semantic Exploration - Fine-tuning VLM
While the main experiments do not use fine-tuned VLMs — since the base performance of the VLM is not a contribution of this work — fine-tuning VLMs with in-domain examples can improve the results and alleviate reliance on performance of existing VLMs. In an additional experiment, we split the 500 scenarios into 300 for training and 200 for testing, and use the 300 training ones for fine-tuning the VLM specifically for question answering. We find, using the same Prismatic VLM, the success rate improves to 68.1% from 56.2%, indicating that fine-tuning on the dataset can better ground VLM’s question answering in the scenario distribution and EQA tasks.
VII-F Q2: Stopping Criterion - Baselines
@@
VII-G Q2: Stopping Criterion - Simulation Results
@@
VII-H Q2: Stopping Criterion - Hardware Results
@@ Acknowledgments
We thank Donovon Jackson, Derick Seale, and Tony Nguyen for contributing to the HM-EQA dataset. The authors were partially supported by the Toyota Research Institute (TRI), the NSF CAREER Award [#2044149], the Office of Naval Research [N00014-23-1-2148], the NSF Award [#1941722], the ONR Award [#N00014-22-1-2293], the DARPA grant [#W911NF2210214], and Princeton SEAS Innovation Award from The Addy Fund for Excellence in Engineering. This article solely reflects the opinions and conclusions of its authors and NSF, ONR, DARPA, TRI or any other Toyota entity. This research was supported in part by Other Transaction award HR00112490375 from the U.S. Defense Advanced Research Projects Agency (DARPA) Friction for Accountability in Conversational Transactions (FACT) program.
@@ Appendix C Additional experiment results
EQA results using other VLMs. The exploration and question answering performance of our method scales with the capabilities of the VLMs. We believe that VLMs have improved significantly and will keep improving over time. For example, there have been new advances in open-source VLMs such as LLaVA 1.6 improving significantly across benchmarks compared to previous ones. We run our exploration experiments with LLaVA 1.6 without fine-tuning, and we found the success rate using the maximum normalized time step improves to 65.3% from 58.4% (Fig. 7 left) achieved in the main experiments. In addition, the official version of state-of-the-art GPT4-V provides access to the log probabilities of its output, which was not available with the preview version. We run our exploration experiments with GPT4-V without fine-tuning, and we found the success rate improves to 73.9% from 58.4%. These results highlight the better question answering and exploration capabilities of the VLMs. Nonetheless, we believe we have not fully exploited the strong semantic reasoning capabilities of the VLMs, and there remains challenges in how to better incorporate, for example, multiple views of the scene for VLM reasoning and question answering in future work.
@@
[…]
Improving semantic exploration by considering view orientation. View orientation can play an important role in determining the semantic value of a location in the scene, e.g., the object on the sofa can be only seen from the front but not from the rear due to occlusion. Our proposed setup in Section IV-C does not consider this aspect; originally we alleviate the issue by restricting the robot to move in smaller steps such that one semantically relevant location might be seen by the robot from multiple view orientations (Section IV-B). In an additional experiment, we save the view orientation from individual observations in the semantic map, and apply bias to such orientation when sampling the next frontier to navigate to. We find this heuristic slightly improves the performance from 58.4% success rate to 59.3%. Future work could add additional dimensions to the semantic map corresponding to discretized view angles similar to affordance maps used in table-top manipulation tasks [zeng2020tossingbot].